OpenStudy (anonymous):

how to convert 5 degree into minutes from this(9.5)degree No.(2) 20.625 degree=20.625*pi/18 how to solve it more

OpenStudy (anonymous):

@kc_kennylau @nincompoop @mathstudent

OpenStudy (kc_kennylau):

1 degree is 60 minute

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

5*60=300

OpenStudy (kc_kennylau):

20.625 degree = 20.625*pi/180 = 165*pi/1440 = 33*pi/288 = 11*pi/96

OpenStudy (kc_kennylau):

And 9.5 degree = 9.5*60 minute = 570 minute

OpenStudy (anonymous):

but answer is not 300 ...

OpenStudy (kc_kennylau):

Sorry I don't understand

OpenStudy (anonymous):

first question is 9.5 degree ...how we change just 5 degree into minutes

OpenStudy (anonymous):

if we multiply 5*60 then we get 300......

OpenStudy (anonymous):

@paki

OpenStudy (mokeira):

20.625 degrees =1237.5 minutes

OpenStudy (paki):

you want to convert 5 degrees into minutes hmm...?

OpenStudy (anonymous):

http://www.rapidtables.com/convert/number/degrees-to-degrees-minutes-seconds.htm in this above link they look here is 30 seconds

OpenStudy (anonymous):

yes 5 dgrees into minutes

OpenStudy (mokeira):

if 1degree=60 minutes 5 degress=5x60=300minutes

OpenStudy (paki):

yeah agree

OpenStudy (mokeira):

@paki am i right?

OpenStudy (paki):

yah

OpenStudy (mokeira):

@ziqbal103 are you converting 5 degrees or 9.5degrees?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

@paki @Mokeira no it is not answer u can use the calculator ...which is in above link and write there 9.5 degree then u will find 30 minutes

OpenStudy (mokeira):

9.5 degress should be equal to 570minutes

OpenStudy (mokeira):

@ziqbal103 in your calculator the give the answer for 0.5 degrees because i see they have don (9.5-9)

OpenStudy (anonymous):

9.5 degrees ...yes i want to know this y they did subtract

OpenStudy (mokeira):

i am also wondering...but that site looks fishy try this one http://www.unitconversion.org/angle/degrees-to-minutes-conversion.html

OpenStudy (anonymous):

my computer is not responding so taking so much time.... in my book answer is 30..but they didn't explain

OpenStudy (mokeira):

wow....how did it get to 30???...why do they subtract in the first place?

OpenStudy (mokeira):

@aryandecoolest @phi @TylerD

OpenStudy (anonymous):

there are 60 minutes in 1degree... so multiplying by 60(5*60) gives 300 not 30 )

OpenStudy (anonymous):

let see what @ganiesh and @phi tell

OpenStudy (mokeira):

o.5 by 60 will give you 30 @aryandecoolest

OpenStudy (mokeira):

@ganeshie8

OpenStudy (anonymous):

yes what @ganeshie8 tell us....

OpenStudy (anonymous):

ummmmm...like to cryyyyyyyyyyyyyyyy...uhunhhhh ...hhhun...hunnnnn..huuu..hahuuu

OpenStudy (anonymous):

now i understand @Mokeira they took .5 *60

OpenStudy (anonymous):

now the next question @Mokeira

OpenStudy (mokeira):

but i dont understand why they did that @ziqbal103

OpenStudy (phi):

Perhaps they want to change 9.5º to degrees, minutes, seconds ? in which case you get 9º 30'

OpenStudy (mokeira):

ooooh...that makes sense

OpenStudy (phi):

the second question(?) 20.625 degree=20.625*pi/180 is translating from degrees to radians, i.e. 20.625º = 11 pi/ 96 rads

OpenStudy (anonymous):

@phi my answer is the same that u done...but in my solved it is 11pi/180

OpenStudy (phi):

20.625 degrees changed to radians is \[ \frac{11 \pi}{96} \]

OpenStudy (anonymous):

thanks for question 1 but @phi 11pi/180 is real....in my book

OpenStudy (anonymous):

@jumperman

OpenStudy (jumperman):

Disturbing my slumber already?

OpenStudy (anonymous):

@Jumperman what say in easy wording plz

OpenStudy (jumperman):

Excuse me?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

tell me that i hve a question that 20.625*pi/180 ...its answer is 11*pi/120 ..how can i get after solving it @Jumperman

OpenStudy (jumperman):

Don't ask me.

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

@phi chewk matsh chat come fass.t.plzzz